# The non-linear dynamics of axion inflation: a detailed lattice study

### v1

@@ VI Discussion
Acknowledgements: We thank R. von Eckardstein, K. Schmitz, O. Sobol, V. Domcke, Y. Ema and S. Sandner for discussion and for kindly providing output data from the gradient expansion formalism. DGF (ORCID $0$-0002-4005-8915) and NL (ORCID $0$-0003-4285-4701) are supported by Generalitat Valenciana grant PROMETEO/2021/083, and by Spanish Ministerio de Ciencia e Innovación grant PID2023-148162NB-C21 and PID2023-148162NB-C22. NL (ORCID $0$-0003-4285-4701) is supported by “Sabor y Origen de la Materia (Flavour and Origin of Matter) CPI-24-295” project. JL (ORCID $0$-0002-1198-3191), AU (ORCID $0$-0002-0238-8390) and JU (ORCID $0$- 0002-4221-2859) acknowledge support from Eusko Jaurlaritza (IT1628-22) and by the PID2021-123703NB-C21 grant funded by MCIN/AEI/10.13039/501100011033/ and by ERDF; “A way of making Europe”. In particular, AU gratefully acknowledges the support from the University of the Basque Country grant (PIF20/151). This work has been possible thanks to the computing infrastructure of the ARINA and Solaris clusters at the University of the Basque Country, UPV/EHU, the Hyperion cluster from the DIPC Supercomputing Center, the FinisTerrae III cluster at Centro de Supercomputación de Galicia (CESGA), the Lluis Vives cluster at the University of Valencia and the Graviton cluster at Instituto de Física Corpuscular (IFIC).
@@ Appendix B Chiral Basis and Projection
$\vec{v}(\hat{\bf k})$ $=$ $\frac{\hat{e}_{3}\times\hat{\bf k}}{|\hat{e}_{3}\times\hat{\bf k}|}=(-\sin\varphi,\cos\varphi,0)\;,$ (92)
$\vec{u}(\hat{\bf k})$ $=$ $\vec{v}\times\hat{\bf k}=(\cos\theta\cos\varphi,\cos\theta\sin\varphi,-\sin\theta)\;.$ (93)
@@
$A_{i}(\textbf{n}+\hat{\imath}/2)\equiv\frac{1}{N^{3}}\sum_{\tilde{\textbf{n}}}e^{-i\frac{2\pi}{N}\tilde{\textbf{n}}(\textbf{n}+\hat{\imath}/2)}A_{i}(\tilde{\textbf{n}})\;,$ (106)
@@ Appendix C Bunch-Davies phase for the initial conditions
We solve the linear regime for the gauge field on a 1D grid for modes well inside the Hubble radius, $k\gg aH$ for different initial conditions. In the top panel of Fig. 19, we show the evolution of $\Delta_{A}^{(+)}$ at intervals of $\Delta\mathcal{N}=1.7$, and in the bottom panel, the evolution of $\Delta_{A}^{(-)}$. The solution using the initial conditions of Eqs. (61)-(62) is depicted in blue, whereas we use green for distorted initial conditions, where we modify the relative complex phase between the gauge field and the electric field to be 00. The distorted initial conditions show oscillatory artifacts in the spectra.
Figure 19: Evolution of the power spectrum of the gauge field for the positive helicity (top) and the negative helicity (bottom), showing a temporal gap of $\Delta\mathcal{N}=1.7$. Different colours correspond to different initial phase differences between the gauge and electric fields: $\pi/2$ in blue and 00 in green. .
@@
$A^{+}(\mathcal{N},\tilde{\bf n})$ $\simeq$ $A^{+}_{\rm{PS}(k)}e^{ik/(aH)}\,,$ (108)
$A^{-}(\mathcal{N},\tilde{\bf n})$ $\simeq$ $A^{-}_{\rm{PS}(k)}e^{ik/(aH)}\;,$ (109)
$E^{+}(\mathcal{N},\tilde{\bf n})$ $\simeq$ $E^{+}_{\rm{PS}(k)}e^{i[k/(aH)-\Theta(k,k_{\rm tr},\epsilon)\pi/2]}\,,$ (110)
$E^{-}(\mathcal{N},\tilde{\bf n})$ $\simeq$ $E^{-}_{\rm{PS}(k)}e^{i[k/(aH)-\pi/2]}\,,$ (111)
where $k=|{\mathbf{k}}|=k_{\rm IR}|\tilde{\bf n}|$, and $A^{\pm}_{\rm{PS}(k)}$ and $E^{\pm}_{\rm{PS}(k)}$ correspond to realizations of the spectra extracted from the mode by mode linear solution at $\mathcal{N}_{\rm switch}$. The function $\Theta(k,k_{\rm tr},\epsilon)$ accounts for the smooth transition of the relative phase in the excited mode, which we model by
$\Theta(k,k_{\rm tr},\epsilon)=\frac{1}{2}\left[\tanh{\left(\epsilon\log{\frac{k}{k_{\rm tr}}}\right)}+1\right]\;,$ (112)
@@ D.2 Intermediate cutoff, $k_{\rm BD}$
Figure 22: Comparison of the evolution of the gauge field power spectrum where the BD vacuum initial condition is set up to $k_{\rm UV}$ (solid turquoise) and where a cutoff has been applied at $k_{\rm BD}$ (dashed black). The top panel shows the full range $[k_{\rm IR},k_{\rm UV}]$, while the bottom panel provides a zoomed-in view of the cutoff region.
The inclusion of an intermediate cutoff scale, $k_{\rm BD}$, in the initialization procedure eliminates the excessive influence of the quantum BD tail of the electromagnetic contributions on the expansion equations and the EOM, see Sec. IV.3. The cutoff is always set so that it provides sufficient support for modes to become excited in the linear regime. In order to verify the validity of this approach, it is necessary to ensure that modes in the range $[k_{\rm BD},k_{\rm UV}]$ grow properly in the non-linear regime, regardless they start from the vacuum solution or from vanishing amplitude.
We show a comparison between different initialisation in Fig. 22, where we include the case with all modes in $[k_{\rm IR},k_{\rm UV}]$ starting in the BD solution (solid lines) and the case where BD is only set in the range $[k_{\rm IR},k_{\rm BD}]$, where $k_{\rm BD}/m=20$ (dashed lines). We simulate starting from $\mathcal{N}_{\rm switch}$ until the end of non-linear inflation. We observe that the cutoff modes grow at the same rate as when starting from the Bunch-Davies vacuum solution and become excited due to the non-linear dynamics. We also show a zoomed-in view of the cutoff region in the lower panel to provide a detailed comparison between both cases.
@@ D.4 Time integrators
where $\langle...\rangle_{\rm V}\equiv\frac{1}{N^{3}}\sum_{\mathbf{n}}(...)$ stands fo volume averaging. $LHS$ and $RHS$ refer to the left- and right-hand sides of Eq. (48) and (46) respectively, and $LHS_{i}=\Delta_{i}^{-}E_{i}$ (considering no sum over repeated indices).

### v2

@@ VI Discussion
Acknowledgements: We thank R. von Eckardstein, K. Schmitz, O. Sobol, V. Domcke, Y. Ema and S. Sandner for discussion and for kindly providing output data from the gradient expansion formalism. DGF (ORCID 0000-0002-4005-8915) and NL (ORCID 0000-0003-4285-4701) are supported by Generalitat Valenciana grant PROMETEO/2021/083, and by Spanish Ministerio de Ciencia e Innovación grant PID2023-148162NB-C21 and PID2023-148162NB-C22. NL (ORCID 0000-0003-4285-4701) is supported by “Sabor y Origen de la Materia (Flavour and Origin of Matter) CPI-24-295” project. JL (ORCID 0000-0002-1198-3191), AU (ORCID 0000-0002-0238-8390) and JU (ORCID 0000- 0002-4221-2859) acknowledge support from Eusko Jaurlaritza (IT1628-22) and by the PID2021-123703NB-C21 grant funded by MCIN/AEI/10.13039/501100011033/ and by ERDF; “A way of making Europe”. In particular, AU gratefully acknowledges the support from the University of the Basque Country grant (PIF20/151). This work has been possible thanks to the computing infrastructure of the ARINA and Solaris clusters at the University of the Basque Country, UPV/EHU, the Hyperion cluster from the DIPC Supercomputing Center, the FinisTerrae III cluster at Centro de Supercomputación de Galicia (CESGA), the Lluis Vives cluster at the University of Valencia and the Graviton cluster at Instituto de Física Corpuscular (IFIC), with funding contributions from EUROPA EXCELENCIA-2022 Grant No. EUR2022-134028 and the European Union NextGenerationEU (PRTR-C17.I01) and Generalitat Valenciana Grant No. ASFAE/2022/020.
@@ Appendix B Chiral Basis and Projection
We choose $\hat{e}_{3}$ as the reference direction to define the orthonormal basis around $\hat{\bf k}={\bf k}/k$. We define a plane orthogonal to $\hat{\bf k}$ with a couple of unitary vectors $(\vec{u},\vec{v})$ (see Fig. 18):
@@
[…]
@@
[…]
@@ Appendix C Bunch-Davies phase for the initial conditions
We solve the linear regime for the gauge field on a 1D grid for modes well inside the Hubble radius, $k\gg aH$ for different initial conditions. In the top panel of Fig. 19, we show the evolution of $\Delta_{A}^{(+)}$ at intervals of $\Delta\mathcal{N}=1.7$, and in the bottom panel, the evolution of $\Delta_{A}^{(-)}$. The solution using the initial conditions of Eqs. (61)-(62) is depicted in blue, whereas we use green for distorted initial conditions, where we modify the relative complex phase between the gauge field and the electric field to be 0. The distorted initial conditions show oscillatory artifacts in the spectra.
Figure 19: Evolution of the power spectrum of the gauge field for the positive helicity (top) and the negative helicity (bottom), showing a temporal gap of $\Delta\mathcal{N}=1.7$. Different colours correspond to different initial phase differences between the gauge and electric fields: $\pi/2$ in blue and 0 in green. .
@@
[…]
@@ D.2 Intermediate cutoff, $k_{\rm BD}$
Figure 22: Comparison of the evolution of the gauge field power spectrum where the BD vacuum initial condition is set up to $k_{\rm UV}$ (solid turquoise) and where a cutoff has been applied at $k_{\rm BD}$ (dashed black). The top panel shows the full range $k_{\rm IR},k_{\rm UV}]$, while the bottom panel provides a zoomed-in view of the cutoff region.
The inclusion of an intermediate cutoff scale, $k_{\rm BD}$, in the initialization procedure eliminates the excessive influence of the quantum BD tail of the electromagnetic contributions on the expansion equations and the EOM, see Sec. IV.3. The cutoff is always set so that it provides sufficient support for modes to become excited in the linear regime. In order to verify the validity of this approach, it is necessary to ensure that modes in the range $k_{\rm BD},k_{\rm UV}]$ grow properly in the non-linear regime, regardless they start from the vacuum solution or from vanishing amplitude.
We show a comparison between different initialisation in Fig. 22, where we include the case with all modes in $k_{\rm IR},k_{\rm UV}]$ starting in the BD solution (solid lines) and the case where BD is only set in the range $k_{\rm IR},k_{\rm BD}]$, where $k_{\rm BD}/m=20$ (dashed lines). We simulate starting from $\mathcal{N}_{\rm switch}$ until the end of non-linear inflation. We observe that the cutoff modes grow at the same rate as when starting from the Bunch-Davies vacuum solution and become excited due to the non-linear dynamics. We also show a zoomed-in view of the cutoff region in the lower panel to provide a detailed comparison between both cases.
@@ D.4 Time integrators
[…]